OpenStudy (anonymous):

Solve this quadratic equation using the quadratic formula. x^2 + 8x - 5 = 0

OpenStudy (anonymous):

\[\frac{ -b \pm \sqrt{ b^2 - 4ac} }{ 2a }\]

OpenStudy (anonymous):

use the quadratic formula. never fails.

OpenStudy (anonymous):

okay I get that part

OpenStudy (anonymous):

yeah so a = 1 b = 8 c = -5

OpenStudy (anonymous):

plug those numbers in for each variable.

OpenStudy (anonymous):

\[\frac{ -8 \pm \sqrt{ 8^2 - 4(1)(-5)} }{ 2(1) }\]

OpenStudy (anonymous):

can you do this simple algebra?

OpenStudy (anonymous):

ya hang on

OpenStudy (anonymous):

I got \[-8\pm \sqrt{84}\]

OpenStudy (anonymous):

ya

OpenStudy (anonymous):

ya

OpenStudy (anonymous):

well there is going to be two different solutions, what did you get?

OpenStudy (anonymous):

idk really im confused

OpenStudy (anonymous):

\[x = \frac{ -8 + \sqrt{84} }{ 2 }\] and \[x = \frac{ -8 - \sqrt{84} }{ 2 }\]

OpenStudy (anonymous):

Simplify the Radicals.

OpenStudy (anonymous):

okay hang on

OpenStudy (anonymous):

1.16515 -17.16515

OpenStudy (anonymous):

yeah but its easier and better if you left it in terms of radicals.

OpenStudy (anonymous):

what

OpenStudy (anonymous):

\[\sqrt{84} = \sqrt{4 \times 3 \times 7} = \sqrt{2^2 \times 3 \times 7} =2 \sqrt{3 \times 7} = 2\sqrt{21}\]

OpenStudy (anonymous):

okay so it will b \[x=2\sqrt{21}\]

OpenStudy (anonymous):

no.

OpenStudy (anonymous):

sad

OpenStudy (anonymous):

\[x = \frac{ -8 + 2\sqrt{21} }{ 2 }\] and \[x = \frac{ -8 - 2\sqrt{21} }{ 2 }\]